# On the Trade-Off between Stability and Fidelity of Gaussian-Smoothed Saliency Maps

### v1

@@ C.3 Other Details
For the calculation of SSIM, we use the one provided by library scikit-image, and set gaussian_weights to True, sigma to 1.5, use_sample_covariance to False, and specify the data_range argument (to 1 for grey images and 255 for RGB images) according to the official suggestions  [^4] https://scikit-image.org/docs/stable/api/skimage.metrics.html#skimage.metrics.structural_similarity to match the implementation of [34]. To calculate top-k mIoU, we set $k=500$ pixels for both saliency maps, and calculate the average IoU for these pixels with top-k values according to the definition.
@@ D.1 Proof of Proposition 1
We begin by introducing Stein’s lemma, as presented in [21]. This lemma is crucial to our overall analysis, as it allows us to express the gradient in expectation in a more manageable form.
@@ D.2 Proof of Theorem 2
D.3 Upper Bound of Stability Error
@@
Next, we quote the results from [27] showing that each SGD step will not make two different parameters much more divergent.
@@
After all these previous results, our next lemma proves the upper bounds for the stability of saliency maps under two different assumptions of the loss function. This proof refers to some analysis in [13].
@@
According to calculations in [13], we can bound

### v2

@@ C.3 Other Details
For the calculation of SSIM, we use the one provided by library scikit-image, and set gaussian_weights to True, sigma to 1.5, use_sample_covariance to False, and specify the data_range argument (to 1 for grey images and 255 for RGB images) according to the official suggestions  [^5] https://scikit-image.org/docs/stable/api/skimage.metrics.html#skimage.metrics.structural_similarity to match the implementation of (Wang et al.,, 2004). To calculate top-k mIoU, we set $k=500$ pixels for both saliency maps, and calculate the average IoU for these pixels with top-k values according to the definition.
@@ D.1 Proof of Proposition 1
We begin by introducing Stein’s lemma, as presented in (Lin et al.,, 2019). This lemma is crucial to our overall analysis, as it allows us to express the gradient in expectation in a more manageable form.
@@ D.2 Proof of Theorem 2
D.3 Upper Bound of Stability Error for Arbitrary Loss Function
@@
Next, we quote the results from (Nesterov,, 2014) showing that each SGD step will not make two different parameters much more divergent.
@@
After all these previous results, our next lemma proves the upper bounds for the stability of saliency maps under two different assumptions of the loss function. This proof refers to some analysis in (Hardt et al.,, 2016).
@@
According to calculations in (Hardt et al.,, 2016), we can bound